find the probability of rolling at least one double 6 in 24 rolls of two six sided dice
Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

This is a very old problem. Goes back to a gambler by the name of Chevalier de Mere in the year 1717.
.
The probability of a double 6 is 1/36
The probability of not getting a double 6 is 35/36
The probability of not getting a double 6 24 times in a row is therefore (35/36)^24.
.
Thus, the probability of at least one double 6 in 24 rolls is 1−(35/36)^24
= 1-0.5085961238691 = 0.4914%
You can round off to however many decimals they told you to, you didn't say

Answer by ikleyn(52873)

.

Hello, @addingup,

        brilliant solution and beautifully presented (!)

        My congrats (!)


I only want to make a correction to your answer.

The probability under the question is   1-0.5085961238691 = 0.4914 = 49.14%.